OpenStudy (anonymous):

Find the radius of Convergence and Interval of Convergence for the given power series

OpenStudy (anonymous):

\[\sum_{n=1}^\infty \frac{2^n}{n}x^n\]

OpenStudy (anonymous):

\[a_n=\frac{2^n}{n}\] \[a_{n+1}=\frac{2^{n+1}}{n+1}\]

OpenStudy (experimentx):

i guess |x| >= 1/2

OpenStudy (anonymous):

\[|(x-0)|\lim_{n-> \infty} |\frac{\frac{2^n}{n}}{\frac{2^{n+1}}{n+1}}|\]

OpenStudy (anonymous):

\[|x-0| \lim_{n->\infty}|\frac{2^n}{n}*\frac{n+1}{2^{n+1}}|\]

OpenStudy (anonymous):

\[|x-0|\lim_{n->\infty}|\frac{2^n}{n}*\frac{n+1}{2^n*2}\]

OpenStudy (anonymous):

i think you're right

OpenStudy (experimentx):